OpenStudy (anonymous):

help?!?!?!?! find the variables not given a1 = 10, r = 3, n = 5

jimthompson5910 (jim_thompson5910):

This looks like a geometric sequence

OpenStudy (anonymous):

it is

jimthompson5910 (jim_thompson5910):

an = a1*r^(n-1) is the general formula

jimthompson5910 (jim_thompson5910):

In this case, a1 = 10 and r = 3, so an = 10*3^(n-1)

jimthompson5910 (jim_thompson5910):

the nth term, an, in this case is the 5th term since n = 5 so a5 = 10*3^(5-1)

jimthompson5910 (jim_thompson5910):

or an = 10*3^(5-1) when n = 5

OpenStudy (anonymous):

High school math or what?

OpenStudy (anonymous):

so,